# Measuring Social Norms of Large Language Models

### v2

@@ Appendix D Summary of Skills
Subject: Language Arts Skill: draw-inferences-from-a-text Description: Read the text. Kari looked under the table. Then she searched the whole room. "You have got to be kidding!" Kari shouted. Grumbling, Kari tore the puzzle apart and put it back in the box. On the top of the box, she crossed out 1000 and wrote 999 above it. Based on the text, which sentence is more likely to be true? Choices: [(a) Kari doesn’t like building puzzles., (b) Kari’s puzzle is missing a piece.] Answer index: 1
@@
Subject: Language Arts Skill: is-the-sentence-simple-or-compound Description: Which is a **compound sentence**? Choices: [(a) Jayden grew too many pumpkins, so he sold the extras at the market., (b) British general John Burgoyne surrendered to American forces at the Battle of Saratoga.] Answer index: 0
@@
Subject: Language Arts Skill: identify-elements-of-poetry Description: Choose the poem that uses **end rhyme**. Choices: [(a) I break a staff. I break the tough branch. I know no light in the woods. I have lost pace with the winds., (b) Sign of a nation, great and strong Toward her people from foreign wrong: Pride and glory and honor,—all Live in the colors to stand or fall.] Answer index: 1
@@
Subject: Language Arts Skill: read-about-business-and-technology Description: Review the text. title: Where Are You? GPS has also become very useful in the field of science. Scientists have figured out ways to use GPS for monitoring earthquakes and volcanoes. GPS can also be used to measure water levels and snowfall. While working with wild animals, some animal researchers also use GPS devices. These scientists can attach small devices onto animals to track the animals’ movements and gather important data about them. Who knows how GPS will help us next? Which is the best summary of the text? Choices: [(a) GPS receivers help people find their way. They work by sending radio signals to satellites in space. Drivers, people enjoying outdoor sports, and emergency workers all use GPS. GPS can also be used for other purposes, like for tracking animals and for science research., (b) "GPS" stands for Global Positioning System. GPS works anywhere on Earth, except deep in caves and underwater. People use GPS every day to drive and play sports. They also use GPS to find lost pets., (c) GPS devices have become very important in our world. People use them for driving and outdoor sports. Scientists use them to monitor earthquakes, volcanoes, and the phases of the moon. GPS can also be used to remind people to walk their pets.] Answer index: 0
@@
Subject: Language Arts Skill: read-about-history Description: Read the text. title: A Whale of a Record It might seem incredible that just one album could help save the lives of so many whales. But remember that Songs of the Humpback Whale was no ordinary record. This one-of-a-kind album revealed secrets about whales that had long been hidden beneath the waves. It allowed these amazing and mysterious creatures to finally share their songs with the human world. What is the text about? Choices: [(a) It is about how musicians recorded an album to encourage people to ban deep-sea whaling., (b) It is about how whales learned to use sounds to communicate after biologists played them recordings., (c) It is about how people changed their mind about recordings of nature after hearing an album of whale songs., (d) It is about how an album of whale recordings helped increase protection for whales around the world.] Answer index: 3
Subject: Language Arts Skill: identify-plurals-singular-possessives-and-plural-possessives Description: What is the word in bold? Are there any **statues** in your town? Choices: [(a) a singular possessive noun, (b) a plural noun] Answer index: 1
Subject: Language Arts Skill: analyze-the-effects-of-figures-of-speech-on-meaning-and-tone Description: Read the text. The figure of speech is shown in bold. Then I knew that even though **nobody in the world** figured that old car had any good purpose, Ob knew there was some real reason to let it sit. And when May died, he figured out what it was. I never saw two people love each other so much. From Cynthia Rylant, Missing May. Copyright 1992 by Cynthia Rylant What does this example of **hyperbole** tell you? Choices: [(a) It suggests that very few people thought the car had a purpose., (b) It suggests that the car travels around the world.] Answer index: 0
Subject: Language Arts Skill: determine-the-meanings-of-greek-and-latin-roots Description: What does the root **tract** mean? Choices: [(a) bend or curve, (b) pull or drag, (c) far away] Answer index: 1
Subject: Language Arts Skill: is-it-a-telling-sentence-or-an-asking-sentence Description: What kind of sentence is this? My father reads to me. Choices: [(a) telling sentence, (b) asking sentence] Answer index: 0
Subject: Language Arts Skill: find-the-vowel-in-the-word Description: Find the vowel. Choices: [(a) o, (b) n] Answer index: 0
@@
Subject: Language Arts Skill: complete-the-sentence-with-the-correct-short-vowel-word Description: Which word makes the most sense in the sentence? The bug is on his ___. Choices: [(a) nap, (b) leg] Answer index: 1
Subject: Language Arts Skill: complete-the-rhyme Description: Complete the rhyme. The men sit in the ___. Choices: [(a) den, (b) back] Answer index: 0
Subject: Language Arts Skill: capitalize-the-first-letter-of-a-sentence Description: Complete the sentence with an uppercase letter. ___ans can go fast. Choices: [(a) V, (b) v] Answer index: 0
Subject: Language Arts Skill: who-what-when-where-or-why Description: Pick the right question word. ___ did your mom go? Choices: [(a) Who, (b) Where] Answer index: 1
Subject: Language Arts Skill: what-am-i Description: Answer the riddle. I am small. What am I? Choices: [(a) a baby, (b) a box] Answer index: 0
Subject: Language Arts Skill: read-questions-with-short-vowel-words Description: Answer the question. Can a wig jog? Choices: [(a) yes, (b) no] Answer index: 1
Subject: Language Arts Skill: choose-the-uppercase-letter-that-matches-review Description: Pick the uppercase letter that matches. x Choices: [(a) X, (b) W] Answer index: 0
Subject: Language Arts Skill: find-the-action-verb Description: Find the verb. A verb shows an action. Choices: [(a) tree, (b) grow] Answer index: 1
Subject: Language Arts Skill: choose-the-lowercase-letter-that-matches-review Description: Pick the lowercase letter that matches. V Choices: [(a) n, (b) v] Answer index: 1
Subject: Language Arts Skill: find-the-complete-sentence Description: Find the complete sentence. Choices: [(a) car, (b) We ate the cake.] Answer index: 1
@@
Subject: Language Arts Skill: which-word-is-not-like-the-others Description: Which word is not like the others? Choices: [(a) big, (b) tall, (c) little, (d) nice] Answer index: 3
@@
Subject: Language Arts Skill: complete-the-sentence-with-the-correct-sight-word-sets-1-2-3 Description: Complete the sentence. I did not want to trick ___. Choices: [(a) has, (b) as, (c) them] Answer index: 2
@@
Subject: Language Arts Skill: complete-the-sentence-with-the-correct-sight-word-sets-4-5-6-7 Description: Complete the sentence. I ate a hot dog ___ went to bed. Choices: [(a) four, (b) more, (c) then] Answer index: 2
Subject: Language Arts Skill: use-the-correct-modal-verb Description: Select the modal verb that best completes the sentence. The coach ___ let Mark join the team if it weren’t so late in the season. Choices: [(a) would, (b) will] Answer index: 0
Subject: Language Arts Skill: use-the-perfect-verb-tenses Description: Complete the sentence with the correct helping verb or verbs. When I walked by Sophia’s desk, I saw that she ___ fallen asleep! Choices: [(a) had, (b) will have] Answer index: 0
Subject: Language Arts Skill: determine-the-meaning-of-idioms-from-context-set-2 Description: What is the meaning of **get one’s feet wet**? Mrs. Tyler is a professional singer who **got her feet wet** by singing in the school choir as a kid. Choices: [(a) to get soaked or messy, (b) to start learning something] Answer index: 1
Subject: Language Arts Skill: choose-between-the-past-tense-and-past-participle Description: Complete the sentence with the correct form of the verb. I swept up crumbs from the cookie Barry had ___. Choices: [(a) eaten, (b) ate] Answer index: 0
Subject: Language Arts Skill: identify-the-meaning-of-idioms-and-adages-set-2 Description: What is the meaning of **if you can’t stand the heat, get out of the kitchen**? Choices: [(a) follow the customs of the place where you are, (b) don’t take on a job or task if it’s too much for you] Answer index: 1
Subject: Language Arts Skill: is-the-sentence-declarative-interrogative-imperative-or-exclamatory Description: What kind of sentence is this? Nora is reading a magazine about skateboarding. Choices: [(a) interrogative, (b) declarative] Answer index: 1
Subject: Language Arts Skill: use-the-progressive-verb-tenses Description: Complete the sentence with the correct helping verb or verbs. "We ___ eating dinner at seven o’clock, so don’t be late," Mrs. Christensen said. Choices: [(a) were, (b) will be] Answer index: 1
Subject: Language Arts Skill: what-does-the-modal-verb-show Description: What does the modal verb show? Select the best answer choice. Bernie and Gary **will** visit Yosemite National Park next summer. Choices: [(a) a request for permission, (b) future action] Answer index: 1
Subject: Language Arts Skill: identify-the-meaning-of-idioms-and-adages-set-1 Description: What is the meaning of **cut corners**? Choices: [(a) to accidentally take the edges off something, (b) to take shortcuts when working] Answer index: 1
Subject: Language Arts Skill: use-thesaurus-entries Description: Look at this thesaurus entry: **abstain** Synonyms: deny, forgo, give up, quit, refrain Which is something you **abstain** from? Choices: [(a) something you almost always do, (b) something you choose not to do] Answer index: 1
Subject: Language Arts Skill: words-with-sub Description: Which might be a **subtopic** within a discussion of the Revolutionary War? Choices: [(a) the Declaration of Independence, (b) U.S. history] Answer index: 0
@@
Subject: Language Arts Skill: words-with-pre Description: What helps **prevent** a cold? Choices: [(a) washing your hands frequently, (b) taking cold medicine] Answer index: 0
Subject: Language Arts Skill: words-with-less Description: Why might you make a **careless** mistake in your work? Choices: [(a) because you aren’t paying attention, (b) because you don’t understand something] Answer index: 0
Subject: Language Arts Skill: words-with-un-dis-in-im-and-non Description: Which is an **infrequent** visitor? Choices: [(a) one you often see, (b) one you rarely see] Answer index: 1
Subject: Language Arts Skill: words-with-re Description: Why might you **reattempt** a challenge? Choices: [(a) because you didn’t succeed the first time, (b) because it’s a new one] Answer index: 0
Subject: Language Arts Skill: use-the-correct-pair-of-correlative-conjunctions Description: Complete the sentence with the correct correlative conjunctions. Kurt’s horse is in ___ the barn ___ the pasture right now. Choices: [(a) not only . . . but also, (b) either . . . or] Answer index: 1
Subject: Language Arts Skill: words-with-ful Description: How would you feel after a **restful** day? Choices: [(a) recharged, (b) exhausted] Answer index: 0
Subject: Language Arts Skill: compound-subjects-and-objects-with-personal-pronouns Description: Select the correct pronoun to complete the sentence. At the garage sale, ___ and Warren hoped to find some good books. Choices: [(a) she, (b) her] Answer index: 0
Subject: Language Arts Skill: determine-the-themes-of-short-stories Description: Read the following story. Paolo walked out on the dock and looked down into the water. "I only like to swim in pools. I’m not swimming in there," he said. "Well, I’m going to," his mom said. "There’s not much else to do here!" Paolo’s little sister, Celia, was standing next to him. "If Paolo’s not going in, I’m not swimming in there either," Celia said. Paolo’s mom looked at Paolo with a raised eyebrow. "That’s fine," she said. "You two don’t have to swim. But remember, there’s no internet here and no TV." Paolo marched back to the cabin and sat in a lawn chair. He crossed his arms and frowned. "This is going to be an awful vacation", he thought. No Wi-Fi, nothing to do. Celia followed him, sitting in a chair and crossing her arms, too. "I’m bored!" Celia said. Meanwhile, their mom had gone into the cabin to change. Soon she emerged, carrying an armful of goggles and towels. "I’m going for a dip. Are you "sure" you don’t want to go?" she asked. Celia glanced at Paolo. "I’m sure," said Paolo. "No way!" added Celia. Paolo sat in the chair and stared across the lake. Celia did too, but soon she grew restless. She got up, walked up and down the dock, and finally sat down, watching her mom. Seeing this, Paolo felt bad. He went inside and put on his bathing suit. Then he walked out to the dock where Celia was sitting. When their mom swam up to them, Paolo apologized. "I’ll swim with you, Mom. I’m sorry I was crabby. The water actually looks nice," he said. Celia looked at Paolo. Then she ran to the cabin to change into her swimsuit, too. Which of the following best describes the main theme or lesson of the story? Choices: [(a) Set a good example for people who look up to you., (b) People today take technology for granted.] Answer index: 0
@@
Subject: Language Arts Skill: words-with-able-and-ible Description: Which item is not a **collectible**? Choices: [(a) an ordinary spiral notebook, (b) a rare comic book] Answer index: 0
@@
Subject: Language Arts Skill: use-the-correct-frequently-confused-word Description: Complete the text with the correct word. Rhianna made a wire sculpture of a tree for her art project. She used steel wool to make ___ leaves. Choices: [(a) its, (b) it’s] Answer index: 0
Subject: Language Arts Skill: words-with-mis Description: Which type of behavior shouldn’t cause **mistrust**? Choices: [(a) suspicious behavior, (b) honest behavior] Answer index: 1
@@
Subject: Language Arts Skill: which-sentence-is-more-formal Description: Which sentence is more formal? Choices: [(a) Visitors are advised that admission to the aircraft museum has been raised by two dollars., (b) FYI, admission to the aircraft museum has been raised by two dollars.] Answer index: 0
@@
Subject: Language Arts Skill: is-the-sentence-simple-compound-or-complex Description: Which is a **simple sentence**? Choices: [(a) Although I would rather visit Italy, I’m excited about going to Panama this summer., (b) The thermometer and the cough syrup are in the medicine cabinet next to the cotton balls.] Answer index: 1
@@
Subject: Language Arts Skill: choose-topic-sentences-for-expository-paragraphs Description: Read the paragraph. Then, add the better topic sentence. ___ In fact, more than half of the planet is covered just by oceans. Earth also has many lakes, rivers, and swamps full of freshwater. But there’s even more freshwater that is frozen in glaciers and icecaps. Add to that all of the water that’s underground or in the air, and you’ve got one wet planet! Choices: [(a) Most of the Earth is covered in water., (b) The highest point on Earth is Mount Everest.] Answer index: 0
Subject: Language Arts Skill: is-the-sentence-a-statement-question-command-or-exclamation Description: What kind of sentence is this? Bobby gave me the best present! Choices: [(a) an exclamation, (b) a statement] Answer index: 0
@@
Subject: Language Arts Skill: does-the-adjective-tell-you-what-kind-or-how-many Description: Look at the adjective in **bold**. Does it tell you "what kind" or "how many"? Maria will need **seven** stamps. Choices: [(a) what kind, (b) how many] Answer index: 1
Subject: Language Arts Skill: determine-the-meanings-of-similes Description: What is the meaning of the simile? Just as a soft breeze rocks the trees, Mrs. Middleton rocked the baby to sleep. Choices: [(a) Mrs. Middleton rocked the baby outside., (b) Mrs. Middleton rocked the baby gently.] Answer index: 1
Subject: Language Arts Skill: is-the-noun-singular-or-plural Description: Is the noun in bold singular or plural? Lamar shouted when his **foot** touched the hot water. Choices: [(a) singular, (b) plural] Answer index: 0
Subject: Language Arts Skill: use-the-correct-article-a-or-an Description: Complete the sentence with the best article. Mrs. Casey keeps the tools in ___ shed. Choices: [(a) an, (b) a] Answer index: 1
Subject: Language Arts Skill: determine-the-themes-of-myths-fables-and-folktales Description: Read the following fable by Aesop. A fox one day spotted a beautiful bunch of ripe grapes hanging from a vine. The grapes seemed ready to burst with juice. The fox’s mouth watered, and he gazed longingly at them. The bunch hung up high, far out of the fox’s reach. He tried to jump and grab it. The first time he jumped, he was still far away. So he walked off a short distance and took a running leap. But he fell short once more. Again and again he tried, but he simply could not reach the grapes. After a long time, he sat down and looked at the grapes in anger. "What a fool I am," he said. "I am wearing myself out, and for what? This is just a bunch of terrible, sour grapes." And away he walked with his nose in the air. What is the main theme or lesson of the fable? Choices: [(a) Sometimes fruit is sour, not sweet., (b) People pretend they don’t want things they can’t have.] Answer index: 1

### v3

@@ Appendix D Summary of Skills
Subject: Language Arts
Skill: draw-inferences-from-a-text
Description: Read the text.
Kari looked under the table. Then she searched the whole room.
"You have got to be kidding!" Kari shouted.
Grumbling, Kari tore the puzzle apart and put it back in the box. On the top of the box, she crossed out 1000 and wrote 999 above it.
Based on the text, which sentence is more likely to be true?
Choices: [(a) Kari doesn’t like building puzzles., (b) Kari’s puzzle is missing a piece.]
Answer index: 1
Subject: Language Arts
Skill: is-the-sentence-simple-or-compound
Description: Which is a **compound sentence**?
@@
Choices: [(a) Jayden grew too many pumpkins, so he sold the extras at the market., (b) British general John Burgoyne surrendered to American forces at the Battle of Saratoga.]
Answer index: 0
Subject: Language Arts
Skill: identify-elements-of-poetry
Description: Choose the poem that uses **end rhyme**.
@@
Choices: [(a) I break a staff. I break the tough branch. I know no light in the woods. I have lost pace with the winds., (b) Sign of a nation, great and strong Toward her people from foreign wrong: Pride and glory and honor,—all Live in the colors to stand or fall.]
Answer index: 1
Subject: Language Arts
Skill: read-about-business-and-technology
Description: Review the text.
title: Where Are You?
@@
GPS has also become very useful in the field of science. Scientists have figured out ways to use GPS for monitoring earthquakes and volcanoes. GPS can also be used to measure water levels and snowfall. While working with wild animals, some animal researchers also use GPS devices. These scientists can attach small devices onto animals to track the animals’ movements and gather important data about them. Who knows how GPS will help us next?
Which is the best summary of the text?
Choices: [(a) GPS receivers help people find their way. They work by sending radio signals to satellites in space. Drivers, people enjoying outdoor sports, and emergency workers all use GPS. GPS can also be used for other purposes, like for tracking animals and for science research., (b) "GPS" stands for Global Positioning System. GPS works anywhere on Earth, except deep in caves and underwater. People use GPS every day to drive and play sports. They also use GPS to find lost pets., (c) GPS devices have become very important in our world. People use them for driving and outdoor sports. Scientists use them to monitor earthquakes, volcanoes, and the phases of the moon. GPS can also be used to remind people to walk their pets.]
Answer index: 0
Subject: Language Arts
Skill: read-about-history
Description: Read the text.
title: A Whale of a Record
@@
It might seem incredible that just one album could help save the lives of so many whales. But remember that Songs of the Humpback Whale was no ordinary record. This one-of-a-kind album revealed secrets about whales that had long been hidden beneath the waves. It allowed these amazing and mysterious creatures to finally share their songs with the human world.
What is the text about?
Choices: [(a) It is about how musicians recorded an album to encourage people to ban deep-sea whaling., (b) It is about how whales learned to use sounds to communicate after biologists played them recordings., (c) It is about how people changed their mind about recordings of nature after hearing an album of whale songs., (d) It is about how an album of whale recordings helped increase protection for whales around the world.]
Answer index: 3
Subject: Language Arts
Skill: identify-plurals-singular-possessives-and-plural-possessives
Description: What is the word in bold?
Are there any **statues** in your town?
Choices: [(a) a singular possessive noun, (b) a plural noun]
Answer index: 1
Subject: Language Arts
Skill: analyze-the-effects-of-figures-of-speech-on-meaning-and-tone
Description: Read the text. The figure of speech is shown in bold.
Then I knew that even though **nobody in the world** figured that old car had any good purpose, Ob knew there was some real reason to let it sit. And when May died, he figured out what it was.
I never saw two people love each other so much.
From Cynthia Rylant, Missing May. Copyright 1992 by Cynthia Rylant
What does this example of **hyperbole** tell you?
Choices: [(a) It suggests that very few people thought the car had a purpose., (b) It suggests that the car travels around the world.]
Answer index: 0
Subject: Language Arts
Skill: determine-the-meanings-of-greek-and-latin-roots
Description: What does the root **tract** mean?
Choices: [(a) bend or curve, (b) pull or drag, (c) far away]
Answer index: 1
Subject: Language Arts
Skill: is-it-a-telling-sentence-or-an-asking-sentence
Description: What kind of sentence is this?
My father reads to me.
Choices: [(a) telling sentence, (b) asking sentence]
Answer index: 0
Subject: Language Arts
Skill: find-the-vowel-in-the-word
Description: Find the vowel.
Choices: [(a) o, (b) n]
Answer index: 0
Subject: Language Arts
@@
Skill: complete-the-sentence-with-the-correct-short-vowel-word
Description: Which word makes the most sense in the sentence?
The bug is on his ___.
Choices: [(a) nap, (b) leg]
Answer index: 1
Subject: Language Arts
Skill: complete-the-rhyme
Description: Complete the rhyme.
The men sit in the ___.
Choices: [(a) den, (b) back]
Answer index: 0
Subject: Language Arts
Skill: capitalize-the-first-letter-of-a-sentence
Description: Complete the sentence with an uppercase letter.
___ans can go fast.
Choices: [(a) V, (b) v]
Answer index: 0
Subject: Language Arts
Skill: who-what-when-where-or-why
Description: Pick the right question word.
___ did your mom go?
Choices: [(a) Who, (b) Where]
Answer index: 1
Subject: Language Arts
Skill: what-am-i
Description: Answer the riddle.
I am small.
What am I?
Choices: [(a) a baby, (b) a box]
Answer index: 0
Subject: Language Arts
Skill: read-questions-with-short-vowel-words
Description: Answer the question.
Can a wig jog?
Choices: [(a) yes, (b) no]
Answer index: 1
Subject: Language Arts
Skill: choose-the-uppercase-letter-that-matches-review
Description: Pick the uppercase letter that matches.
x
Choices: [(a) X, (b) W]
Answer index: 0
Subject: Language Arts
Skill: find-the-action-verb
Description: Find the verb. A verb shows an action.
Choices: [(a) tree, (b) grow]
Answer index: 1
Subject: Language Arts
Skill: choose-the-lowercase-letter-that-matches-review
Description: Pick the lowercase letter that matches.
V
Choices: [(a) n, (b) v]
Answer index: 1
Subject: Language Arts
Skill: find-the-complete-sentence
Description: Find the complete sentence.
Choices: [(a) car, (b) We ate the cake.]
Answer index: 1
@@
Subject: Language Arts
Skill: which-word-is-not-like-the-others
Description: Which word is not like the others?
Choices: [(a) big, (b) tall, (c) little, (d) nice]
Answer index: 3
Subject: Language Arts
@@
Skill: complete-the-sentence-with-the-correct-sight-word-sets-1-2-3
Description: Complete the sentence.
I did not want to trick ___.
Choices: [(a) has, (b) as, (c) them]
Answer index: 2
Subject: Language Arts
@@
Skill: complete-the-sentence-with-the-correct-sight-word-sets-4-5-6-7
Description: Complete the sentence.
I ate a hot dog ___ went to bed.
Choices: [(a) four, (b) more, (c) then]
Answer index: 2
Subject: Language Arts
Skill: use-the-correct-modal-verb
Description: Select the modal verb that best completes the sentence.
The coach ___ let Mark join the team if it weren’t so late in the season.
Choices: [(a) would, (b) will]
Answer index: 0
Subject: Language Arts
Skill: use-the-perfect-verb-tenses
Description: Complete the sentence with the correct helping verb or verbs.
When I walked by Sophia’s desk, I saw that she ___ fallen asleep!
Choices: [(a) had, (b) will have]
Answer index: 0
Subject: Language Arts
Skill: determine-the-meaning-of-idioms-from-context-set-2
Description: What is the meaning of **get one’s feet wet**?
Mrs. Tyler is a professional singer who **got her feet wet** by singing in the school choir as a kid.
Choices: [(a) to get soaked or messy, (b) to start learning something]
Answer index: 1
Subject: Language Arts
Skill: choose-between-the-past-tense-and-past-participle
Description: Complete the sentence with the correct form of the verb.
I swept up crumbs from the cookie Barry had ___.
Choices: [(a) eaten, (b) ate]
Answer index: 0
Subject: Language Arts
Skill: identify-the-meaning-of-idioms-and-adages-set-2
Description: What is the meaning of **if you can’t stand the heat, get out of the kitchen**?
Choices: [(a) follow the customs of the place where you are, (b) don’t take on a job or task if it’s too much for you]
Answer index: 1
Subject: Language Arts
Skill: is-the-sentence-declarative-interrogative-imperative-or-exclamatory
Description: What kind of sentence is this?
Nora is reading a magazine about skateboarding.
Choices: [(a) interrogative, (b) declarative]
Answer index: 1
Subject: Language Arts
Skill: use-the-progressive-verb-tenses
Description: Complete the sentence with the correct helping verb or verbs.
"We ___ eating dinner at seven o’clock, so don’t be late," Mrs. Christensen said.
Choices: [(a) were, (b) will be]
Answer index: 1
Subject: Language Arts
Skill: what-does-the-modal-verb-show
Description: What does the modal verb show? Select the best answer choice.
Bernie and Gary **will** visit Yosemite National Park next summer.
Choices: [(a) a request for permission, (b) future action]
Answer index: 1
Subject: Language Arts
Skill: identify-the-meaning-of-idioms-and-adages-set-1
Description: What is the meaning of **cut corners**?
Choices: [(a) to accidentally take the edges off something, (b) to take shortcuts when working]
Answer index: 1
Subject: Language Arts
Skill: use-thesaurus-entries
Description: Look at this thesaurus entry:
*abstain**
Synonyms: deny, forgo, give up, quit, refrain
Which is something you **abstain** from?
Choices: [(a) something you almost always do, (b) something you choose not to do]
Answer index: 1
Subject: Language Arts
Skill: words-with-sub
Description: Which might be a **subtopic** within a discussion of the Revolutionary War?
Choices: [(a) the Declaration of Independence, (b) U.S. history]
Answer index: 0
Subject: Language Arts
Skill: words-with-pre
Description: What helps **prevent** a cold?
@@
Choices: [(a) washing your hands frequently, (b) taking cold medicine]
Answer index: 0
Subject: Language Arts
Skill: words-with-less
Description: Why might you make a **careless** mistake in your work?
Choices: [(a) because you aren’t paying attention, (b) because you don’t understand something]
Answer index: 0
Subject: Language Arts
Skill: words-with-un-dis-in-im-and-non
Description: Which is an **infrequent** visitor?
Choices: [(a) one you often see, (b) one you rarely see]
Answer index: 1
Subject: Language Arts
Skill: words-with-re
Description: Why might you **reattempt** a challenge?
Choices: [(a) because you didn’t succeed the first time, (b) because it’s a new one]
Answer index: 0
Subject: Language Arts
Skill: use-the-correct-pair-of-correlative-conjunctions
Description: Complete the sentence with the correct correlative conjunctions.
Kurt’s horse is in ___ the barn ___ the pasture right now.
Choices: [(a) not only . . . but also, (b) either . . . or]
Answer index: 1
Subject: Language Arts
Skill: words-with-ful
Description: How would you feel after a **restful** day?
Choices: [(a) recharged, (b) exhausted]
Answer index: 0
Subject: Language Arts
Skill: compound-subjects-and-objects-with-personal-pronouns
Description: Select the correct pronoun to complete the sentence.
At the garage sale, ___ and Warren hoped to find some good books.
Choices: [(a) she, (b) her]
Answer index: 0
Subject: Language Arts
Skill: determine-the-themes-of-short-stories
Description: Read the following story.
Paolo walked out on the dock and looked down into the water. "I only like to swim in pools. I’m not swimming in there," he said.
"Well, I’m going to," his mom said. "There’s not much else to do here!"
Paolo’s little sister, Celia, was standing next to him. "If Paolo’s not going in, I’m not swimming in there either," Celia said.
Paolo’s mom looked at Paolo with a raised eyebrow. "That’s fine," she said. "You two don’t have to swim. But remember, there’s no internet here and no TV."
Paolo marched back to the cabin and sat in a lawn chair. He crossed his arms and frowned. "This is going to be an awful vacation", he thought. No Wi-Fi, nothing to do.
Celia followed him, sitting in a chair and crossing her arms, too. "I’m bored!" Celia said.
Meanwhile, their mom had gone into the cabin to change. Soon she emerged, carrying an armful of goggles and towels. "I’m going for a dip. Are you "sure" you don’t want to go?" she asked.
Celia glanced at Paolo.
"I’m sure," said Paolo.
"No way!" added Celia.
Paolo sat in the chair and stared across the lake. Celia did too, but soon she grew restless. She got up, walked up and down the dock, and finally sat down, watching her mom.
Seeing this, Paolo felt bad. He went inside and put on his bathing suit. Then he walked out to the dock where Celia was sitting. When their mom swam up to them, Paolo apologized. "I’ll swim with you, Mom. I’m sorry I was crabby. The water actually looks nice," he said.
Celia looked at Paolo. Then she ran to the cabin to change into her swimsuit, too.
Which of the following best describes the main theme or lesson of the story?
Choices: [(a) Set a good example for people who look up to you., (b) People today take technology for granted.]
Answer index: 0
@@
Subject: Language Arts
Skill: words-with-able-and-ible
Description: Which item is not a **collectible**?
Choices: [(a) an ordinary spiral notebook, (b) a rare comic book]
Answer index: 0
Subject: Language Arts
Skill: use-the-correct-frequently-confused-word
Description: Complete the text with the correct word.
@@
Rhianna made a wire sculpture of a tree for her art project. She used steel wool to make ___ leaves.
Choices: [(a) its, (b) it’s]
Answer index: 0
Subject: Language Arts
Skill: words-with-mis
Description: Which type of behavior shouldn’t cause **mistrust**?
Choices: [(a) suspicious behavior, (b) honest behavior]
Answer index: 1
Subject: Language Arts
Skill: which-sentence-is-more-formal
Description: Which sentence is more formal?
@@
Choices: [(a) Visitors are advised that admission to the aircraft museum has been raised by two dollars., (b) FYI, admission to the aircraft museum has been raised by two dollars.]
Answer index: 0
Subject: Language Arts
Skill: is-the-sentence-simple-compound-or-complex
Description: Which is a **simple sentence**?
@@
Choices: [(a) Although I would rather visit Italy, I’m excited about going to Panama this summer., (b) The thermometer and the cough syrup are in the medicine cabinet next to the cotton balls.]
Answer index: 1
Subject: Language Arts
Skill: choose-topic-sentences-for-expository-paragraphs
Description: Read the paragraph. Then, add the better topic sentence.
@@
___ In fact, more than half of the planet is covered just by oceans. Earth also has many lakes, rivers, and swamps full of freshwater. But there’s even more freshwater that is frozen in glaciers and icecaps. Add to that all of the water that’s underground or in the air, and you’ve got one wet planet!
Choices: [(a) Most of the Earth is covered in water., (b) The highest point on Earth is Mount Everest.]
Answer index: 0
Subject: Language Arts
Skill: is-the-sentence-a-statement-question-command-or-exclamation
Description: What kind of sentence is this?
Bobby gave me the best present!
Choices: [(a) an exclamation, (b) a statement]
Answer index: 0
Subject: Language Arts
Skill: does-the-adjective-tell-you-what-kind-or-how-many
@@
Description: Look at the adjective in **bold**. Does it tell you "what kind" or "how many"?
Maria will need **seven** stamps.
Choices: [(a) what kind, (b) how many]
Answer index: 1
Subject: Language Arts
Skill: determine-the-meanings-of-similes
Description: What is the meaning of the simile?
Just as a soft breeze rocks the trees, Mrs. Middleton rocked the baby to sleep.
Choices: [(a) Mrs. Middleton rocked the baby outside., (b) Mrs. Middleton rocked the baby gently.]
Answer index: 1
Subject: Language Arts
Skill: is-the-noun-singular-or-plural
Description: Is the noun in bold singular or plural?
Lamar shouted when his **foot** touched the hot water.
Choices: [(a) singular, (b) plural]
Answer index: 0
Subject: Language Arts
Skill: use-the-correct-article-a-or-an
Description: Complete the sentence with the best article.
Mrs. Casey keeps the tools in ___ shed.
Choices: [(a) an, (b) a]
Answer index: 1
Subject: Language Arts
Skill: determine-the-themes-of-myths-fables-and-folktales
Description: Read the following fable by Aesop.
A fox one day spotted a beautiful bunch of ripe grapes hanging from a vine. The grapes seemed ready to burst with juice. The fox’s mouth watered, and he gazed longingly at them.
The bunch hung up high, far out of the fox’s reach. He tried to jump and grab it. The first time he jumped, he was still far away. So he walked off a short distance and took a running leap. But he fell short once more. Again and again he tried, but he simply could not reach the grapes.
After a long time, he sat down and looked at the grapes in anger.
"What a fool I am," he said. "I am wearing myself out, and for what? This is just a bunch of terrible, sour grapes."
And away he walked with his nose in the air.
What is the main theme or lesson of the fable?
Choices: [(a) Sometimes fruit is sour, not sweet., (b) People pretend they don’t want things they can’t have.]
Answer index: 1
Subject: Language Arts
Skill: to-be-use-the-correct-form